# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ 9. $f^{\delta}_{\mathrm{RR}}$-concurrent composition
Lemma (Restatement of Lemma 5.9).
@@
We begin by proving some intermediate results needed for the final proof. Vadhan et al. [24] show that to guarantee an IM is differentially private, it suffices to consider only deterministic adversaries. Specifically, the following holds:
Lemma 9.1 ([24]).
An IM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. a neighboring relation $\sim$ if and and only if for every deterministic adversary $\mathcal{A}$ and every pair of neighboring initial states $s\sim s^{\prime}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s^{\prime}))$ are $(\epsilon,\delta)$-indistinguishable.
@@
A CM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$ if and and only if for every deterministic adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
@@
Let $Y^{0}$ and $Y^{1}$ be random variables with domain $\mathcal{Y}$. For every (not necessarily maximal) coupling $(Z^{0},Z^{1})$ of $(Y^{0},Y^{1})$, where the marginals of $Z^{0}$ and $Z^{1}$ are identically distributed as of $Y^{0}$ and $Y^{1}$, we have
@@
  • If the histories are identical, i.e. $Z_{1:i-1}^{0}=Z_{1:i-1}^{1}=y_{1:i-1}$, then $(Z_{i}^{0},Z_{i}^{1})$ is drawn from the maximal coupling of $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}$}, given by Lemma 9.6.
@@
A backward iteration with the same proof implies that $\Pr[Z^{0}=Z^{1}]\geq 1-\gamma$. Consequently, $\Pr[Z^{0}\neq Z^{1}]\leq\gamma$, and by Lemma 9.7, $\mathrm{TV}(Y^{0},Y^{1})\leq\gamma$. ∎
@@
  • If $w_{i-1}=\mathdollar$ (i.e., the communication has halted before) or the design of $\mathcal{A}$ is such that it halts the communication after receiving the i1i-1-st response $w_{i-1}$ given the history $w_{1:i-1}$, then $W_{1:i-1}=\mathdollar$ with probability $1$.
@@
Define $\Delta_{i}(w_{1:i-1})=0$ in the former case and $\Delta_{i}(w_{1:i-1})=\delta_{i}^{\prime}(w_{1:i-1},\mathcal{A})$ in the latter. By Lemma 9.4 and Lemma 2.4, for every $i\in\mathbb{N}$ and every $w_{1:i-1}\in S_{i-1}^{b}$,
@@ 10. Concurrent Filter Composition
Given a filter $\mathrm{Filt}$, we define a verification function $f[\mathrm{Filt}]$ as follows and refer to the $f[\mathrm{Filt}]$-concurrent composition of CMs as the concurrent $\mathrm{Filt}$-filter composition of CMs. The verification function $f[\mathrm{Filt}]$ verifies the validity of message formats, extracts the privacy parameters from the creation queries, and apply $\mathrm{Filt}$ on them.
@@
Rogers et al. [21] studied the $\mathrm{Filt}$-filter composition of NIMs for the first time. More recently, Haney et al. [11] showed that if the $\mathrm{Filt}$-filter composition of NIMs is $(\epsilon,\delta)$-DP, then the same guarantee extends to the $\mathrm{Filt}$-filter composition of IMs. We further generalize this result to the $\mathrm{Filt}$-filter composition of CMs. To state our result, we need to define the $\mathrm{Filt}$-filter composition of randomized response mechanisms:
@@
For every filter $\mathrm{Filt}$, if the $\mathrm{Filt}$-filter composition of randomized response mechanisms is $(\epsilon,\delta)$-DP, then the $\mathrm{Filt}$-filter composition of CMs is also $(\epsilon,\delta)$-DP.
@@
Let $f=f^{\mathrm{Filt}}$ and $f_{\mathrm{RR}}=f^{\mathrm{Filt}}_{\mathrm{RR}}$. This proof is identical to the proof of Theorem 4.9, except the following modifications: First, the right mechanism of the IPM $\mathcal{P}$ is $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ instead of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$. Second, the variable $\sigma_{1}$ is set differently. In Theorem 4.9, $\mathcal{P}$ initially interacted with its right mechanism, ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$, and set $\sigma_{1}$ to its output sequence $(r_{1},\dots,r_{k})$, without ever changing it. Here, $\mathcal{P}$ interacts with its right mechanism, $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, each time it receives a pair of creation queries as a left message and appends a randomized response $r_{i}$ to $\sigma_{1}$.
Specifically, the same as before, the state of $\mathcal{P}$ consists of two sequences $\sigma_{1}$ and $\sigma_{2}$, both initialized empty. Upon receiving a left message of the form $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})^{2}$, $\mathcal{P}$ takes the following steps: First, $\mathcal{P}$ appends the pair $(\mathcal{T}_{i},t_{i})$ to the sequence $\sigma_{2}$, where $\mathcal{T}_{i}$ is the IPM corresponding to $\mathcal{M}_{i}$ from Lemma 4.6 and $t_{i}$ is its initial state. Then, it sends the message $(\beta_{\epsilon_{i},\delta_{i}},\beta_{\epsilon_{i},\delta_{i}})$ to its right mechanism $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, where $\beta_{\epsilon_{i},\delta_{i}}$ is the creation query corresponding to $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$. Upon receiving the acknowledgment response $\top$ from $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, $\mathcal{P}$ discards it and sends the message $\left((0,i),(1,i)\right)$ to its right mechanism again. This message will be transformed by $\mathcal{I}(b)$ to $(b,i)$. Receiving $(b,i)$, the mechanism ExtConComp sends $b$ to $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ and returns its response $r_{i}$, which will be forwarded unchanged to $\mathcal{P}$. $\mathcal{P}$ then appends $r_{i}$ to $\sigma_{1}$. Finally, $\mathcal{P}$ returns the acknowledgment message $\top$ to its left mechanism.

### v4

@@ 11. $f^{\delta}_{\mathrm{RR}}$-concurrent composition
Lemma 0 (Restatement of Lemma 5.9).
@@
We begin by proving some intermediate results needed for the final proof. Vadhan et al. (vadhan2022concurrent, ) show that to guarantee an IM is differentially private, it suffices to consider only deterministic adversaries. Specifically, the following holds:
Lemma 11.1 ((vadhan2022concurrent, )).
An IM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. a neighboring relation $\sim$ if and only if for every deterministic adversary $\mathcal{A}$ and every pair of neighboring initial states $s\sim s^{\prime}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s^{\prime}))$ are $(\epsilon,\delta)$-indistinguishable.
@@
A CM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$ if and only if for every deterministic adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
@@
Let $Y^{0}$ and $Y^{1}$ be random variables with domain $\mathcal{Y}$. For every (not necessarily maximal) coupling $(Z^{0},Z^{1})$ of $Y^{0}$ and $Y^{1}$, where the marginals $Z^{0}$ and $Z^{1}$ are distributed as $Y^{0}$ and $Y^{1}$, we have
@@
  • If the histories are identical, i.e. $Z_{1:i-1}^{0}=Z_{1:i-1}^{1}=y_{1:i-1}$, then $(Z_{i}^{0},Z_{i}^{1})$ is drawn from the maximal coupling of teh distributions of $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}$}, given by Lemma 11.6.
@@
A backward iteration with the same proof shows that $\Pr[Z^{0}=Z^{1}]\geq 1-\gamma$. Consequently, $\Pr[Z^{0}\neq Z^{1}]\leq\gamma$, and by Lemma 11.7, $\mathrm{TV}(Y^{0},Y^{1})\leq\gamma$. ∎
@@
  • If $w_{i-1}=\mathdollar$ (i.e., the communication has halted before) or the design of $\mathcal{A}$ is such that it halts the communication after receiving the (i1)(i-1)-th response $w_{i-1}$ given the history $w_{1:i-1}$, then $W_{1:i-1}=\mathdollar$ with probability $1$.
@@
Define $\Delta_{i}(w_{1:i-1})=0$ in the former case and $\Delta_{i}(w_{1:i-1})=\delta_{i}^{\prime}(w_{1:i-1},\mathcal{A})$ in the latter. By Lemmas 11.4 and 2.4, for every $i\in\mathbb{N}$ and every $w_{1:i-1}\in S_{i-1}^{b}$,
@@ 12. Concurrent Filter Composition
Given a filter $\mathrm{Filt}$, we define a verification function $f^{\mathrm{Filt}}$ as follows and refer to the $f^{\mathrm{Filt}}$-concurrent composition of CMs as the concurrent $\mathrm{Filt}$-filter composition of CMs. The verification function $f^{\mathrm{Filt}}$ verifies the validity of message formats, extracts the privacy parameters from the creation queries, and applies $\mathrm{Filt}$ on them.
@@
Rogers et al. (rogers2016privacy, ) studied the $\mathrm{Filt}$-filter composition of NIMs for the first time. More recently, Haney et al. (haney2023concurrent, ) showed that if the $\mathrm{Filt}$-filter composition of NIMs is $(\epsilon,\delta)$-DP, then the same guarantee extends to the concurrent $\mathrm{Filt}$-filter composition of IMs. We further generalize this result to the concurrent $\mathrm{Filt}$-filter composition of CMs. To state our result, we need to define the $\mathrm{Filt}$-filter composition of randomized response mechanisms:
@@
For every filter $\mathrm{Filt}$, if the $\mathrm{Filt}$-filter composition of randomized response mechanisms is $(\epsilon,\delta)$-DP, then the concurrent $\mathrm{Filt}$-filter composition of CMs is also $(\epsilon,\delta)$-DP.
@@
Let $f=f^{\mathrm{Filt}}$ and $f_{\mathrm{RR}}=f^{\mathrm{Filt}}_{\mathrm{RR}}$. This proof is identical to the proof of Theorem 4.10, except for the following modifications: First, the right mechanism of the IPM $\mathcal{P}$ is $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ instead of ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$. Second, the variable $\sigma_{1}$ is set differently. In Theorem 4.10, $\mathcal{P}$ initially interacted with its right mechanism, ${\textrm{Comp}}[\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}]\left((b)_{i=1}^{k}\right)$, and set $\sigma_{1}$ to its output sequence $(r_{1},\dots,r_{k})$, without ever changing it. Here, $\mathcal{P}$ interacts with its right mechanism, $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, each time it receives a pair of creation queries as a left message and appends a randomized response $r_{i}$ to $\sigma_{1}$.
Specifically, as before, the state of $\mathcal{P}$ consists of two sequences $\sigma_{1}$ and $\sigma_{2}$, both initialized empty. Upon receiving a left message of the form $(\mathcal{M}_{i},s_{i},\epsilon_{i},\delta_{i},f_{i})^{2}$, $\mathcal{P}$ takes the following steps: First, $\mathcal{P}$ appends the pair $(\mathcal{T}_{i},t_{i})$ to the sequence $\sigma_{2}$, where $\mathcal{T}_{i}$ is the IPM corresponding to $\mathcal{M}_{i}$ from Lemma 4.7 and $t_{i}$ is its initial state. Then, it sends the message $(\beta_{\epsilon_{i},\delta_{i}},\beta_{\epsilon_{i},\delta_{i}})$ to its right mechanism $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, where $\beta_{\epsilon_{i},\delta_{i}}$ is the creation query corresponding to $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$. Upon receiving the acknowledgment response $\top$ from $\mathcal{V}[f_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$, $\mathcal{P}$ discards it and sends the message $\left((0,i),(1,i)\right)$ to its right mechanism again. This message will be transformed by $\mathcal{I}(b)$ to $(b,i)$. Receiving $(b,i)$, the mechanism ExtConComp sends $b$ to $\mathrm{RR}_{\epsilon_{i},\delta_{i}}$ and returns its response $r_{i}$, which will be forwarded unchanged to $\mathcal{P}$. $\mathcal{P}$ then appends $r_{i}$ to $\sigma_{1}$. Finally, $\mathcal{P}$ returns the acknowledgment message $\top$ to its left mechanism.